A commercial company uses capital of 70 million UM., of which 50% represents the fixed capital and pays salaries of 25 million UM. The fixed capital is depreciated in 7 years. Production value is 70 million UM.

Determine the annual cost of production, the profit mass and the profit rate calculated on turnover.


1
Expert's answer
2022-02-14T15:26:46-0500

"Annual Cost Of Production=Variable Capital+Salaries"

"Variable Capital=(70000000)-(\frac{50}{100}×70000000)"

"=35000000"

"Annual Cost=35000000+25000000=60000000"

Profit mass"=Production Value-AnnualCost"

"=70000000-60000000=10000000"

Profit Rate"=\frac{10000000}{60000000}×100=16.7" %